OpenStudy (anonymous):

(3(d^-1)-5)-2/(3(d^-4)-2 answer in positive exponents step by step please

OpenStudy (anonymous):

i think i might have did something wrong

OpenStudy (anonymous):

alright well let me check if this is what you mean \[\frac{3d^{-1}-2}{3d^{-4}-2}\]

OpenStudy (anonymous):

if so when you have negatives like that ou can look at it like \[\frac{3d^4-2}{3d^1-2}\]

OpenStudy (anonymous):

umm not exactly its like this \[3(d ^{-1})^{-5})^{-2}/(3(d ^{-4})^{-2}\]

OpenStudy (anonymous):

sorry for the bottom part its \[(3(d ^{-4})^{-2})^{-2}\]

OpenStudy (anonymous):

and the 3 is within the parenthesis/

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

the top will come out to be \[3(d^{-1*-5*-2})\] i believe as your syntax is off by a bit i beileve

OpenStudy (anonymous):

unless that 3 is in the parenthesis n the top also

OpenStudy (anonymous):

it is lol

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

\[3^{-2}d^8\]

OpenStudy (anonymous):

i believe will be the bottom

OpenStudy (anonymous):

no wait... ^-16

OpenStudy (anonymous):

geez it wont let me see the problems

OpenStudy (anonymous):

i'm getting \[\frac{3^{-2}d^{-10}}{3^{-2}d^{-16}}\]

OpenStudy (anonymous):

when you multiply those out

OpenStudy (anonymous):

\[=\frac{\frac{1}{9}d^{-10}}{\frac{1}{9}d^{-16}}\]

OpenStudy (anonymous):

1/9 will cancel

OpenStudy (anonymous):

flip them and then subtract = d^6

OpenStudy (anonymous):

i got 1/d6 iono if thats right

OpenStudy (anonymous):

upload your work

OpenStudy (anonymous):

oo i kno why its because i forgot to include that the number can be negative its just that the exponents cant be negative

OpenStudy (anonymous):

o nvm your right i did have a mistake

OpenStudy (anonymous):

if you upload your work i can maybe see where you got off track...

OpenStudy (anonymous):

i canceled the -6 but i forgot to flip the negative exponents and then subtracted them which gave me another answer i went back and looked over it and saw where i had the mistake

OpenStudy (anonymous):

ah yes , if you plan on going into calculus you will do this a lot, so i'd work more on the whole manipulation of exponents, as it will help when you get there =]

OpenStudy (anonymous):

you can also cancel everything and it will leave you d^6 i have done this before its just that it has been awhile and i forgot everything lol and yea thanks :]